How much of an alloy that is 40% copper should be mixed with 100 ounces of an alloy that is 60% copper in order to get an alloy that is 50% copper?
Answer by ewatrrr(24785)

 
Hi,
Let x represent how much of an alloy that is 40% copper that is needed:
.40x + .60*100 = .50(100+x)
.10*100 = .10x
.10*100/.10 = x
100 = x
CHECKING our Answer***
.40*100 + .60*100 = 40 + 60 = 100 = .50*200